OpenStudy (anonymous):

help please Question 36 Solve. 2x2 + x = 15. A) x = 2 and x = –15 B) x = 17 and x = 13 C) x = and x = –3 D) x = –2 and x =

OpenStudy (anonymous):

for d) its x=5/3

OpenStudy (anonymous):

according to this question x can have only ONE and only ONE answer. if x has two answers, x shouls be squared? is it tha case?

OpenStudy (anonymous):

im assuming..

OpenStudy (mertsj):

\[2x^2+x-15=0\] \[(2x-5)(x+3)=0\] \[x=\frac{5}{2}, -3\]

OpenStudy (anonymous):

(2x-5)(x+3) x=-3 or 5/2

OpenStudy (mertsj):

Choice C.

OpenStudy (anonymous):

thnaks so much :)

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

the above answers are right. :) but heiress when typing a power use ^ :)

OpenStudy (anonymous):

that's y i was confused

OpenStudy (anonymous):

k :) sorry